Square root of 18 + Square root of 12.
I dont have a square root button on my computer..sadly, so hopefully you see what im asking.

Answer by stanbon(75887)

Square root of 18 + Square root of 12.
sqrt(18)=sqrt(9*2)=3sqrt2
sqrt(12)=sqrt(4*2)=2sqrt2
Adding them you get 5sqrt2.
That is probably the answer you wnat.
You don't really need a square-root button.
Cheers,
Stan H.